When the digits of a two-digit number are reversed, the new number is 9 more than the original number, and the sum of the digits of the original number is 11. What is the original number?
Answer by stanbon(75887)

When the digits of a two-digit number are reversed, the new number is 9 more than the original number, and the sum of the digits of the original number is 11. What is the original number?
---------------
Let the original number be 10t+u where t is the tens digit and u is the units
digit.
-----------------
When reversed the number would be 10u+t
------------------
EQUATION:
10u+t=10t+u+9
t+u=11
---------
Rearrange to get:
9u-9t=9
u-t=1
u+t=11
----------
Add the last two to solve for "u":
2u = 12
u = 6
Substitute to solve for "t":
6+t=11
t = 5
-----------
Original Number: 56
========================
Cheers,
Stan H.